Two electronic speakers with 250Hz and 260Hz frequency are set on. Determine the beat frequency and the perceived frequency heard by the observer.


1
Expert's answer
2021-02-18T18:56:16-0500

Beat Frequency


"f_B=260-250=10\ Hz"

Perceived Frequency


"f_o=0.5(260+250)=255\ Hz"